OpenStudy (anonymous):

Approximate the real number solution(s) to the polynomial function f(x) = x3 – 4x2 + x + 6 . x = –3, x = –1, x = 2 x = 3, x = 1, x = –2 x = 3, x = –1, x = 2 x = –3, x = 1, x = –2

OpenStudy (asnaseer):

what methods have you been told to use? also, I assume the question is actually: Approximate the real number solution(s) to the polynomial function f(x) = x3 – 4x2 + x + 6 = 0

OpenStudy (asnaseer):

i.e. find the x-intercepts

OpenStudy (anonymous):

thats all it says

OpenStudy (asnaseer):

what methods have you been taught in class?

OpenStudy (asnaseer):

e.g. you could approximate the solutions by drawing a graph of f(x) and checking where it crosses the x axes

OpenStudy (anonymous):

how would that work?

OpenStudy (asnaseer):

what do you mean by "work"?

OpenStudy (asnaseer):

if you just have:\[f(x) = x^3 – 4x^2 + x + 6\]then this has "no solutions" - it is just defining a curve given by the polynomial. However, if you were asked to find where this curve crosses on the x axes, then effectively you are finding solutions to:\[0=x^3 – 4x^2 + x + 6\]and one of doing that is to draw a graph of this curve and see where it crosses the x axes.

OpenStudy (anonymous):

would it be B?

OpenStudy (asnaseer):

have you drawn the curve?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

can you upload a picture of it here please so I can check where you may have made a mistake?

OpenStudy (anonymous):

oh wait, C